When gamma rays are incident on matter, the intensity of the gamma rays passing through the material varies with depth x as I(x) = Ioe^(-ux) , where Io is the intensity of the radiation at the surface of the material (at x = 0) and u is the linear absorption coefficient. For 0.400 MeV gamma rays in lead, the linear absorption coefficient is 1.59cm^-1 .
(a) Determine the "half-thickness" for lead, that is, the thickness of lead that would absorb half the incident gamma rays.
(b) What thickness reduces the radiation by a factor of 10^4?
1
Expert's answer
2021-02-23T10:06:55-0500

a) According to the definition of intensity, we have the formula for the intensity of radiation as

"I = \frac{E}{At}"

From this formula, intensity I is directly proportional to the energy E of the radiation.

When gamma rays are incident on matter, the intensity of the gamma rays passing through the material varies with depth x as

"I(x) = I_0e^{-\u03bcx}"

Where "I_0" is the intensity of the radiation at the surface of the material, and μ is the absorption coefficient for lead.

"I_0 = 0.400 \;MeV \\

I = \frac{I_0}{2} \\

I(x) = I_0e^{-μx} \\

\frac{I_0}{2} = I_0e^{-μx} \\

\frac{1}{2} = e^{-μx} \\

= e^{-(1.59 \;cm^{-1})x} \\

-(1.59 \;cm^{-1})x = ln(\frac{1}{2}) \\

= -0.69315 \\

x = \frac{-0.69315}{-1.59 \;cm^{-1}} \\

= 0.44 \;cm"

b) It is given that

"I = \frac{I_0}{10^4} \\

= 10^{-4}I_0 \\

I(x) = I_0e^{-μx} \\

10^{-4}I_0 = e^{-(1.59 \;cm^{-1})x} \\

-(1.59 \;cm^{-1} x) = ln(10^{-4}) \\

= -9.21034 \\

x =\frac{-9.21034}{-1.59 \;cm^{-1}} \\

= 5.793 \;cm"